OpenStudy (anonymous):

10x+7y= 49 10y-x=70 use either elimination or substitution

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

u familiar wid substitution method... ?

OpenStudy (anonymous):

idk this is a review from algebra I, im in geometry, so i don't remember how to do algebra stuff

ganeshie8 (ganeshie8):

its okay :) its simple, il walk you through

OpenStudy (anonymous):

ok thanks

ganeshie8 (ganeshie8):

Given equations : 10x+7y= 49 ------(1) 10y-x=70 --------(2)

ganeshie8 (ganeshie8):

from (2), 10y-x=70 x = 10y - 70, substitute this x value in (1)

ganeshie8 (ganeshie8):

10x+7y= 49 10(10y-70) + 7y = 49 100y - 700 + 7y = 49 107y = 749 y = ?

OpenStudy (anonymous):

7. ok thanks!

ganeshie8 (ganeshie8):

y=7, you need to find x also

OpenStudy (anonymous):

aw mann

OpenStudy (anonymous):

so how

OpenStudy (anonymous):

plug it in the 1st equation?

ganeshie8 (ganeshie8):

yes, looks you're back to algebra brilliance again :)

OpenStudy (anonymous):

haha ok

OpenStudy (anonymous):

i sitll have 11 more oh mygod!

OpenStudy (anonymous):

i got 10x=0 so its no solution right?

OpenStudy (anonymous):

yea it is, i remember my teacher saying that one is no solution

ganeshie8 (ganeshie8):

10x = 0 x = 0

ganeshie8 (ganeshie8):

so the solution of the system is, x = 0 y = 7

OpenStudy (anonymous):

yup